A voltaic cell employs the following redox reaction:
2Fe3+(aq)+3Mg(s)→2Fe(s)+3Mg2+(aq)2Fe3+(aq)+3Mg(s)→2Fe(s)+3Mg2+(aq)
Calculate the cell potential at 25 ∘C∘C under each of the following conditions.

Part A:

standard conditions

Part B:

[Fe3+]=[Fe3+]= 2.0×10−3 MM ; [Mg2+]=[Mg2+]= 2.25 M

Part C:

[Fe3+]=[Fe3+]= 2.25 MM ; [Mg2+]=[Mg2+]= 2.0×10−3 M
Answer #1

A)
from data table:
Eo(Mg2+/Mg(s)) = -2.372 V
Eo(Fe3+/Fe(s)) = -0.04 V

As per given reaction/cell notation,
cathode is (Fe3+/Fe(s))
anode is (Mg2+/Mg(s))

Eocell = Eocathode - Eoanode
= (-0.04) - (-2.372)
= 2.332 V
Answer: 2.33 V

B)
Number of electron being transferred in balanced reaction is 6
So, n = 6

use:
E = Eo - (2.303*RT/nF) log {[Mg2+]^3/[Fe3+]^2}

Here:
2.303*R*T/F
= 2.303*8.314*298.0/96500
= 0.0591

So, above expression becomes:
E = Eo - (0.0591/n) log {[Mg2+]^3/[Fe3+]^2}
E = 2.332 - (0.0591/6) log (2.25^3/0.002^2)
E = 2.332-(6.361*10^-2)
E = 2.268 V
Answer: 2.27 V

C)

Number of electron being transferred in balanced reaction is 6
So, n = 6

use:
E = Eo - (2.303*RT/nF) log {[Mg2+]^3/[Fe3+]^2}

Here:
2.303*R*T/F
= 2.303*8.314*298.0/96500
= 0.0591

So, above expression becomes:
E = Eo - (0.0591/n) log {[Mg2+]^3/[Fe3+]^2}
E = 2.332 - (0.0591/6) log (0.002^3/2.25^2)
E = 2.332-(-8.673*10^-2)
E = 2.419 V
Answer: 2.42 V